# Modelling the interaction between ethnicity and infectious disease transmission dynamics

### v1

@@ 4.2 Limitations
We also treat the first Omicron wave as a SEIR epidemic that ignores differences in time-dependent case rates, which are not well explained by a SEIR model. We fitted the model to data on reported cases with an assumption about how these relate to actual infections. We have investigated a scenario with lower CAR for Māori and Pacific Peoples, but this still assumes some measure of relationship between each ethnicity’s CAR, which may or may not be representative of the real world. This could potentially be partly addressed in future work by including hospitalisations and deaths in the model, which are less sensitive to test-seeking behaviour (though they are sensitive to other differences, e.g. in rates of comorbidity).
@@
Pacific Peoples, Asians, and Europeans/Others are underrepresented in prioritised datasets. For example, when the Ministry of Social Development swapped from a prioritised measure to a total measure, it saw the population identifying as Pacific and European rise from 9% to 12% and 39% to 53%, respectively [53]. Pacific Peoples are also historically underrepresented in the best available population data [54]. Both of these can lead to errors in estimates of model parameters, which are difficult to account for.
@@ Acknowledgements
We are grateful to Kevin C. Ma and Stephen Kissler for their useful discussion about contact matrices.
@@
This work was supported by the Ngāi Tahu Research Centre, Canterbury University and Te Niwha Infectious Diseases Research Platform, Institute of Environmental Science and Research, grant number TN/P/24/UoC/MP and the Marsden Fund, grant number 24-UOC-020.
Appendix A Statistical area information
@@
Table A contains various statistics about the statistical area datasets used in parameter fitting.
[…]
@@
Table of various statistical measures of the ethnic populations within the SA datasets, where SD is the standard deviation of the populations, and interquartile range is the range between the 75th and 25th quartile of the population sizes. Total Mean Median SD Range Interquartile range SA1 Total measure Māori 775410 26.5 18.0 28.2 0 - 435 9 - 33 Pacific Peoples 381576 13.0 6.0 25.4 0 - 315 0 - 12 Asians 707499 24.2 12.0 33.4 0 - 789 3 - 33 Europeans/Others 3424572 117.0 114.0 49.8 0 - 825 84 - 147 All ethnicities 5289057 180.7 174.0 60.4 27 - 1188 141 - 213 SA2 Total measure Māori 775875 362.6 252.0 355.7 0 - 3477 138 - 472 Pacific Peoples 381711 178.4 60.0 381.4 0 - 3501 24 - 138 Asians 707541 330.6 126.0 467.6 0 - 4530 45 - 414 Europeans/Others 3425964 1600.9 1539.0 827.3 9 - 4230 995 - 2136 All ethnicities 5291091 2472.5 2395.5 1243.7 33 - 6099 1518 - 3319 SA2 Priority measure Māori 888840 412.1 290.0 396.6 0 - 3780 160 - 540 Pacific Peoples 359480 166.7 50.0 376.7 0 - 3440 20 - 120 Asians 820580 380.4 150.0 534.0 0 - 5760 50 - 470 Europeans/Others 3048470 1413.3 1330.0 821.0 10 - 7600 830 - 1910 All ethnicities 5117370 2372.4 2300.0 1215.5 10 - 10120 1470 - 3170
Appendix B Parameter Variation
@@
In Figure 7, the transmission rates were held constant to the values fit to the proportionate mixing scheme while $\epsilon$ varied. This was done without the effect of vaccines considered (i.e. vaccinated population as set to zero). As can be seen, the disparity in attack rate between groups widened as $\epsilon$ increased. This was expected, as when epsilon approached one, different ethnicities’ equations decoupled into separate SEIR models, which had no dependence on other ethnicities transmission rates. $\epsilon$ was also varied while the transmission rates were refit to the attack rate data (under the assumed value for CAR). It would not be qualitatively reasonable if the gap between the times of the peak positions was too large and values over 0.6 tend to not have qualitatively reasonable gaps in peak timing.
@@
Appendix C Results for total measure SA1 and SA2 data
Values of the transmission rates fit to the SA1 and SA2 total measure population projections can be seen in Table C.
[…]
Table of transmission rate values fit using the assortative mixing model for total population data. $a_{k}$ CAR $\epsilon$ Māori Pacific Asian European/Other SA1 4060 0.163 0.577 0.929 0.252 0.284 40 0.185 0.519 0.836 0.359 0.408 50 0.192 0.465 0.682 0.340 0.382 60 0.195 0.436 0.617 0.328 0.367 SA2 4060 0.123 0.580 0.947 0.246 0.278 40 0.152 0.521 0.846 0.357 0.406 50 0.158 0.466 0.692 0.337 0.380 60 0.161 0.438 0.626 0.325 0.365
Figures 8 compare the assortative transmission matrices fit using the total measure population data. Figures 9 and 10 show quantification analysis results for the total measure statistical area datasets.
@@ Appendix C Results for total measure SA1 and SA2 data
Figures 9 and 10 show quantification analysis repeated for the total population measures.
Figure 9: Quantification analysis using SA1 total measure population data. The four scenarios considered are: 1) ethnicity-specific parameters and assortativity, 2) no assortativity, 3) population-averaged vaccination rates, 4) population-averaged transmission rates.
Figure 10: Quantification analysis using SA2 total measure population data. The four scenarios considered are: 1) ethnicity-specific parameters and assortativity, 2) no assortativity, 3) population-averaged vaccination rates, 4) population-averaged transmission rates.

### v2

@@ 5.6.2 Model fitting and parameters
Ideally we would have seroprevalence data to complement reported cases to reduce bias present in the data; unfortunately, seroprevalence results were not available for the Omicron wave and the only seroprevalence study in Aotearoa was from earlier in the pandemic [carlton_charting_2021]. We fitted the model to data on reported cases and tested several different assumptions about how these relate to actual infections. We have investigated a scenario with lower CAR for Māori and Pacific Peoples, but this still assumes some measure of relationship between each ethnicity’s CAR, which may or may not be representative of the real world. This could potentially be partly addressed in future work by including hospitalisations and deaths in the model, which are less sensitive to test-seeking behaviour (though they are sensitive to other differences, e.g. in rates of comorbidity).
5.6.3 Ethnic groupings
@@
All considered ethnic groups, excluding Māori, are underrepresented in prioritised datasets. For example, when the Ministry of Social Development swapped from a prioritised measure to a total measure, it saw the population identifying as Pacific and European rise from 9% to 12% and 39% to 53%, respectively [ministry_of_social_development_improving_2021]. Pacific Peoples are also historically underrepresented in the best available population data [sonder_selective_2024]. Both of these can lead to errors in estimates of model parameters, which are difficult to account for.
5.7 Conclusion
@@ Acknowledgements
The authors are grateful to Kevin C. Ma and Stephen Kissler for their useful discussion about contact matrices, to Andrew Sporle, Samik Datta and Nicole Satherley for discussions about modelling infectious disease transmission in different ethnicity groups, and to three anonymous peer reviewers for comments on an earlier version of this manuscript. MJP would like to thank the Banff International Research Station for Mathematical Innovation and Discovery and the organizers and participants of the workshop Mathematical and Statistical Challenges in Post-Pandemic Epidemiology and Public Health (25w5369) held at Casa Matemática Oaxaca in Mexico.
@@
The project “Improving models for pandemic preparedness and response: modelling differences in infectious disease dynamics and impact by ethnicity” (TN/P/24/UoC/MP) was funded by Te Niwha, the Infectious Diseases Research Platform – co-hosted by PHF Science and the University of Otago and provisioned by the Ministry of Business, Innovation and Employment, New Zealand. This research was supported by the Marsden Fund grant (24-UOC-020) managed by Royal Society Te Apārangi. VL would like to acknowledge support from the Ngāi Tahu Research Centre, University of Canterbury.
6 Appendix - supplementary
6.1 Statistical area information
@@
Table 1 contains various summary statistics about the statistical area datasets used in parameter fitting.
Total Mean Median SD Range Interquartile
range
SA1 Total measure
Māori 775410 26.5 18.0 28.2 0 - 435 9 - 33
Pacific Peoples 381576 13.0 6.0 25.4 0 - 315 0 - 12
Asians 707499 24.2 12.0 33.4 0 - 789 3 - 33
Europeans/Others 3424572 117.0 114.0 49.8 0 - 825 84 - 147
All ethnicities 5289057 180.7 174.0 60.4 27 - 1188 141 - 213
SA2 Total measure
Māori 775875 362.6 252.0 355.7 0 - 3477 138 - 472
Pacific Peoples 381711 178.4 60.0 381.4 0 - 3501 24 - 138
Asians 707541 330.6 126.0 467.6 0 - 4530 45 - 414
Europeans/Others 3425964 1600.9 1539.0 827.3 9 - 4230 995 - 2136
All ethnicities 5291091 2472.5 2395.5 1243.7 33 - 6099 1518 - 3319
SA2 Priority measure
Māori 888840 412.1 290.0 396.6 0 - 3780 160 - 540
Pacific Peoples 359480 166.7 50.0 376.7 0 - 3440 20 - 120
Asians 820580 380.4 150.0 534.0 0 - 5760 50 - 470
Europeans/Others 3048470 1413.3 1330.0 821.0 10 - 7600 830 - 1910
All ethnicities 5117370 2372.4 2300.0 1215.5 10 - 10120 1470 - 3170
@@
Table 1: Table of various statistical measures of the ethnic populations within the SA datasets, where SD is the standard deviation of the populations, and interquartile range is the range between the 75th and 25th quartile of the population sizes.
6.2 Transmission rate values
Values of the transmission rates fit to the SA1 and SA2 total measure population projections can be seen in Table 2.
$a_{k}$
CAR $\epsilon$ Māori Pacific Asian European/Other
SA1
40/60 0.162 0.617 1.019 0.281 0.328
40 0.188 0.556 0.916 0.403 0.476
50 0.195 0.496 0.742 0.380 0.444
60 0.198 0.465 0.668 0.366 0.425
SA2
40/60 0.122 0.622 1.039 0.275 0.321
40 0.155 0.558 0.928 0.400 0.473
50 0.160 0.498 0.753 0.377 0.441
60 0.163 0.467 0.678 0.363 0.423
Table 2: Table of transmission rate values fit using the assortative mixing model for total population data.
6.3 Column summation property
The columns of our matrix must sum to the transmission rates (described by the below equation).
$\sum_{i}C_{ij}=a_{j}$
We refer to this as the column summation condition. This condition requires the sum of contacts an individual has with each ethnic group equals the contact rate of the individual’s ethnicity.
6.3.1 Frequency-based assumption
The frequency-based empirically estimated contact matrix is defined via the following equation,
$T_{ij}^{\prime}$ $=\sum_{l}\frac{a_{i}a_{j}N_{i,l}N_{j,l}}{\sum_{k}a_{k}N_{k,l}}$ (20)
This leads to the following definition of the transmission matrix
$C_{ij,\,\text{frequency}}$ $=\frac{1}{N_{j}}\sum_{l}\frac{a_{i}a_{j}N_{i,l}N_{j,l}}{\sum_{k}a_{k}N_{k,l}}$
where the index $l$ represents a block and $N_{kl}$ represents an ethnicity’s population within that block.
This matrix does conserve the meaning of the transmission rates as seen by the following derivation
$\sum_{i}C_{ij,\,\text{frequency}}$ $=\sum_{i}\frac{1}{N_{j}}\sum_{l}\frac{a_{i}a_{j}N_{i,l}N_{j,l}}{\sum_{k}a_{k}N_{k,l}}$
$=\frac{1}{N_{j}}\sum_{l}\frac{a_{j}N_{j,l}\sum_{i}a_{i}N_{i,l}}{\sum_{k}a_{k}N_{k,l}}$
$=\frac{1}{N_{j}}\sum_{l}a_{j}N_{j,l}$
$=a_{j}$
6.3.2 Density-based assumption
The density-based empirically estimated contact matrix is defined by Ma et al. via the following equation,
$\beta_{ij,\,\text{density}}$ $=\frac{\sum_{k}a_{k}N_{k}}{\sum_{i,j,l}a_{i}a_{j}N_{i,l}N_{j,l}}\sum_{l}a_{i}a_{j}N_{i,l}N_{j,l}$
This matrix does not satisfy the column summation condition.
6.4 Parameter Variation
@@
The transmission rates were held constant to the values fit to the proportionate mixing scheme while $\epsilon$ varied. As can be seen in Figure 6, the disparity in attack rate between groups widened as $\epsilon$ increased. This was expected, as when epsilon approached one, different ethnicities’ equations decoupled into separate SEIR models, which had no dependence on other ethnicities transmission rates. $\epsilon$ was also varied while the transmission rates were refit to the attack rate data (under the assumed value for CAR). It would not be qualitatively reasonable if the gap between the times of the peak positions was too large and values over 0.6 tend to not have qualitatively reasonable gaps in peak timing.
@@
6.5 Additional plots
Maps of the ethnic composition of each SA2 region in Aotearoa can be seen in Figure 7. There is high ethnic clustering of the Pacific and Asian groups in cities, particularly Auckland (Figure 8.
Figure 7: Proportion of each SA2 that is a) Māori, b) Pacific, c) Asian, and d) European/Other.
Figure 8: Proportion of each SA2 that is a) Māori, b) Pacific, c) Asian, and d) European/Other in the majority of Auckland SA2 regions.
Figures 9 compare the assortative transmission matrices fit using the total measure population data. Figure 10 shows quantification analysis repeated for total measure statistical area population data.
@@ 6.5 Additional plots
Figure 10: Quantification analysis using a) the SA1 and b) the SA2 total measure population data. The four scenarios considered are: 1) ethnicity-specific parameters and assortativity, 2) no assortativity, 3) population-averaged vaccination rates, 4) population-averaged transmission rates.